The average number of inches of rain in Tampa in August is one inch more than three times the average rainfall in December. The combined rainfall in August and December is 9.8 inches, find the rainfall for each month.
Answer by checkley71(8403)

9.8=(1+3d)+d
9.8=1+4d
4d=9.8-1
4d=8.8
d=8.8/4
d=2.2 inches for December
2.2+1=3.2 inches for August.
proof
9.8=1+3*2.2+3.2
9.8=1+6.6+2.2
9.8=9.8